# Exactly factorized molecular Kohn–Sham density functional theory

### v1

@@ II.1 Exact factorization of molecular wavefunctions
As readily seen from Eq. (14), the conditional Eq. (13) is not a regular electronic equation, also because of the geometrical derivatives. As pointed out in Ref. 43 (see also the references therein) and further discussed in Sec. V, second-order derivatives (last term on the right-hand side of Eq. (14)) prevent us from describing exactly the conditional electronic structure with a single Slater determinant, even when electrons do not interact among themselves, like in a KS molecule Fromager and Lasorne (2024). In order to highlight these additional complications (when comparison is made with regular BO electronic structure theory) and possibly suggest how to deal with them (see Secs. III.2 and V), we can rewrite Eq. (13) in the following matrix form,
@@
$\hat{H}^{(0)}(R)=\hat{H}^{\rm BO}(R)+{U}^{\rm coup(0)}_{\rm ne}(R),$ (18)
and the conditional wavefunction has been replaced by the two-component “wavefunction and its geometrical derivative” one,
thus reducing to first order only the geometrical derivatives present in the Hamiltonian (matrix). Interestingly, the non-Hermitian structure of the latter matrix echoes that of $\hat{U}^{\rm coup[2]}_{\rm ne}$ in the electronic Hilbert space or, more precisely, of its first-order geometrical derivative contribution (second term on the right-hand side of Eq. (14)). This feature, which has already been discussed in the literature Li et al. (2018, 2025), becomes even more apparent when Eq. (17) is rewritten equivalently as follows,
@@ II.2 Molecular KS-DFT
respectively Fromager and Lasorne (2024). Note that, for practical purposes, standard (purely electronic) density-functional approximations (DFAs) can be recycled within the adiabatic approximation that was deduced from an exact molecular generalization of the adiabatic connection formalism in Ref. Fromager and Lasorne (2024). This showed that the molecular Hxc functional can be written as [see Sec. IV of Ref Fromager and Lasorne (2024)]:
$\mathcal{E}_{\rm Hxc}[\Gamma,n]=\int dR\;\Gamma(R)\,\int_{0}^{1}d\lambda\,\langle\hat{W}_{ee}\rangle_{\phi_{R}^{\lambda}[\Gamma,n]},$ (31)
with $\lambda$ being the electron-electron interaction strength scaling parameter and $\phi_{R}^{\lambda}[\Gamma,n]({\bf r})=\frac{\Psi^{\lambda}[\Gamma,n]({R,\bf r})}{\Gamma(R)}$ the effective electronic wavefunction of the partially-interacting molecule. Assuming that, for all $R$, $\phi_{R}^{\lambda}[\Gamma,n]({\bf r})$ matches its local-in-$R$ approximation $\phi^{\lambda}[n_{R}]({\bf r})$, that is the ground partially-interacting BO electronic state reproducing the electronic density at each nuclear geometry, one can leverage the KS-DFT adiabatic connection formula
$E_{\rm Hxc}[n]=\int_{0}^{1}\langle\hat{W}_{ee}\rangle_{\phi^{\lambda}[n]},$ (32)
@@
In this approximate picture, the electronic density $n_{R}:\mathbf{r}\mapsto n(R,\mathbf{r})$ is thought as being parameterized by the geometry $R$, hence the name “adiabatic” given to the approximation. Its limitations and remedies will be discussed in a forthcoming paper (see also Refs. 23; 44; 43). Later in Sec. IV, the theory will be applied to a model system for which numerically exact densities can be evaluated and the corresponding KS potentials can be determined by reverse-engineering.
@@ III.2 Approximate KS-like one-electron conditional equation
thus isolating second-order geometrical derivatives in the second term on the right-hand side. The existence of the latter term echoes the usual considerations related to the account (or not) of diagonal BO corrections (DBOCs) and reduced mass corrections within a BH picture Scherrer et al. (2017); Mátyus and Teufel (2019); Maskri and Joubert-Doriol (2022), which are known to be usually small (perturbative) around equilibrium geometries (far from crossings) and often neglected within the BO approximation. While it might be treated within perturbation theory Li et al. (2025); Tu and Gross (2025b); Schild et al. (2016), the above matrix form indicating clearly and explicitly how the (to-be-defined) unperturbed conditional solutions would couple, a more involved diagonalization approach can also be formulated, as shown later in Sec. V. In this section, we will simply neglect second-order geometrical derivatives, thus leading to the approximate conditional KS equation,
@@
where the index $\mu\geq 0$ has been introduced for labelling a specific solution. The ground-state ($\mu=0$) solution is an approximation to the conditional KS electronic wavefunction $\Phi^{\rm KS}_{R}(r)$. As discussed in Sec. V, excited-state ($\mu>0$) solutions to this problem can be used as a basis for approaching $\Phi^{\rm KS}_{R}(r)$ even further. In the context of perturbation theory, they would play the role of the perturbers. Let us remark that the Hamiltonian is not Hermitian, and its eigenstates are not orthogonal. However the eigenvalues are real-valued.
@@
according to the definition of $\hat{h}^{(0)}(R)$ (see Eqs. (37) and (39d)). Note that, in practice, the zeroth-order coupling potential ${u}^{\rm coup(0)}_{\rm ne}(R)$ would be evaluated from the ground-state ($\mu=0$) solution via the KS vector potential expression (see Eq. (38c)) combined with Eqs. (15) and (39b).
@@
Let us finally stress that Eq. (50) is more advanced than the beyond-BO KS Eq. (40) of Ref. 14, because of the additional first-order geometrical derivative contribution ${u}^{\rm coup(1)}_{\rm ne}(R)\partial/\partial R$. It is also substantially different from the exactly-factorized DFT of Wang et al. Wang et al. (2025) (see also Ref. 43), where the beyond-BO KS potential operator remains multiplicative (see Eq. (17) in Ref. 44).
@@ IV Model system study
The BO ground and first excited electronic PESs are drawn on top panel of Fig 1 together with the exact ground state nuclear density as a function of $R$. The latter was obtained by the means of a Discrete Variable Representation (DVR) Colbert and Miller (1992) calculation of the molecular ground-state wavefunction. On the bottom panel, we compare the electronic density difference between atomic sites obtained in the BO approximation to the exact result. The BO solution undergoes a sharp charge transfer at the avoided crossing between $E_{1}$ and $E_{2}$, going from a strongly ionic character to an almost symmetrical split of electronic density on each atom. In contrast, the exact solution shows taking non-adiabatic effects into account displaces the distance of equilibrium between ionic and neutral character by 1 atomic unit and makes the transition much smoother. The model presents strong non-adiabatic effects while being simple enough so that we can reverse-engineer the KS potentials reproducing the exact (beyond-BO) nuclear and electronic densities. We leave the detailed exposition of our parametrization and numerical details to a follow-up publication, as our present goal is only to compare the prediction of Eq. (48) to a full account of second-order geometric derivatives (see Eq. (40)) for given nuclear-nuclear and nuclear-electron (simply referred to as electronic in the following) KS potentials.
@@
On Fig. 2 we plot the geometric first derivative of the exact KS coefficients in the singlet basis ($\Phi_{R}^{\rm KS}=\sum^{3}_{i=1}C^{\rm KS}_{i}(R)\;\Phi_{i}$). Those were obtained by solving the KS analog to molecular Schrödinger Eq. (1) in full (full lines), with exact nuclear density and electronic KS potential as input. As a comparison, we computed the prediction that Eq. (48) [which is equivalent to the one-electron conditional KS Eq. (50)] gives for the coefficients’ first derivative when fed with exact KS coefficients, electronic potential and nuclear density (dashed lines). The singular behavior around $R=4$ a.u. corresponds to the maximum of the ground-state nuclear density and thus simply reflects that when ${u}^{\rm coup(1)}_{\rm ne}(R)$ vanishes, Eq. (48) does not tell us any information about the change of the electronic wavefunction at this position. There, it simplifies to a (local) eigenvalue problem with a KS electronic potential imbuing non-local (beyond-BO) effects. Thus, would the approximate solution be satisfactory everywhere else, no pathologic behavior would actually emerge. On that point, the prediction of Eq. (48) is reasonably close to the exact one so that our proposed decomposition into a ”first order” solution (Eq. (50)) followed by a perturbative incorporation of second-order derivative contributions holds promise.
@@ V Outlook: Correlating the conditional KS wavefunction
$[{\bf S}(R)]_{\nu\mu}=\int dr\,\left[\Phi^{\nu}_{R}(r)\right]^{*}\Phi^{\mu}_{R}(r),$ (60)
respectively. The non-orthogonality of the Slater determinants stems from the non-Hermitian nature of the “first-order” approximate Hamiltonian in Eq. (49). In general, such an operator has complex eigenvalues, but cases where they happen to be real exist Surján et al. (2024), as in the present situation, due to the nature of the terms involved in assembling their expressions. A working equation can finally be obtained through multiplication by the marginal nuclear density and integration over the geometry, thus leading to
@@
As a final note, let us point out that a possibly better starting point for correlating the conditional KS wavefunction would consist in incorporating second-order geometrical derivatives into the orbitals as follows,
@@
This would generate a new set of Slater determinants
${\Phi}^{\mu}_{R}\rightarrow\tilde{\Phi}^{\mu}_{R}{\equiv}\left|\tilde{\varphi}_{R}^{(1^{\mu})}\tilde{\varphi}_{R}^{(2^{\mu})}\ldots\tilde{\varphi}_{R}^{(N_{\rm e}^{\mu})}\right|$ (67)
that fulfill the more accurate (but still approximate) conditional KS equation,
@@
where $\frac{\partial^{(1)}}{\partial R}$ operates onto the first-order Slater determinant geometrical derivative $\partial\tilde{\Phi}^{\mu}_{R}(r)/\partial R$ like a one-electron operator, i.e.,
$\begin{split}&\frac{\partial^{(1)}}{\partial R}\dfrac{\partial\tilde{\Phi}^{\mu}_{R}}{\partial R}\\ &\equiv\sum^{N_{\rm e}}_{i=1}\left|\tilde{\varphi}_{R}^{(1^{\mu})}\ldots\dfrac{\partial^{2}\tilde{\varphi}_{R}^{(i^{\mu})}}{\partial R^{2}}\ldots\tilde{\varphi}_{R}^{(N_{\rm e}^{\mu})}\right|,\end{split}$ (69)
and
@@
On that basis, the correlated conditional KS wavefunction would be expanded as follows
@@
and the non-orthogonal CI problem rewritten from the following separation of the conditional KS Hamiltonian matrix,
$\begin{split}&\left[\begin{matrix}\hat{h}^{(0)}(R)&{u}^{\rm coup(1)}_{\rm ne}(R)-\frac{1}{2M}\frac{\partial}{\partial R}\\ \frac{\partial}{\partial R}&0\end{matrix}\right]\\ &=\left[\begin{matrix}\hat{h}^{(0)}(R)&{u}^{\rm coup(1)}_{\rm ne}(R)-\frac{1}{2M}\frac{\partial^{(1)}}{\partial R}\\ \frac{\partial}{\partial R}&0\end{matrix}\right]\\ &\quad+\left[\begin{matrix}0&-\frac{1}{2M}\frac{\partial^{(2)}}{\partial R}\\ 0&0\end{matrix}\right],\end{split}$ (72)
where
$\frac{\partial^{(2)}}{\partial R}\frac{\partial\tilde{\Phi}^{\mu}_{R}(r)}{\partial R}=\dfrac{\partial^{2}\tilde{\Phi}^{\mu}_{R}(r)}{\partial R^{2}}-\frac{\partial^{(1)}}{\partial R}\dfrac{\partial\tilde{\Phi}^{\mu}_{R}}{\partial R}$ (73)
describes the first-order geometrical derivative of two different orbitals within the Slater determinant $\tilde{\Phi}^{\mu}_{R}$ and does model, as such, correlation effects.

### v2

@@ II.1 Exact factorization of molecular wavefunctions
As readily seen from Eq. (14), the conditional Eq. (13) is not a regular electronic equation, also because of the geometrical derivatives. As pointed out in Ref. 43 (see also the references therein) and further discussed in Sec. V, second-order derivatives (last term on the right-hand side of Eq. (14)) prevent us from describing exactly the conditional electronic structure with a single Slater determinant, even when electrons do not interact among themselves, like in a KS molecule Fromager and Lasorne (2024). In order to highlight these additional complications (when comparison is made with regular BO electronic structure theory) and possibly suggest how to deal with them (see Secs. III.2 and V), it is instructive, by analogy with second-order differential equations in time of classical mechanics, for example, to rewrite Eq. (13) as a first-order equation in $R$ instead. This can be achieved by using, as unknown quantity, the conditional wavefunction and its geometrical derivative, together,
thus leading to the equivalent matrix form of Eq. (13),
@@
$\hat{H}^{(0)}(R)=\hat{H}^{\rm BO}(R)+{U}^{\rm coup(0)}_{\rm ne}(R).$ (19)
Interestingly, the non-Hermitian structure of the latter matrix echoes that of $\hat{U}^{\rm coup[2]}_{\rm ne}$ (see Eq. (14)) in the electronic Hilbert space or, more precisely, of its first-order geometrical derivative contribution (second term on the right-hand side of Eq. (14)). This feature, which has already been discussed in the literature Li et al. (2018, 2025), becomes even more apparent when Eq. (18) is rewritten equivalently as follows,
@@
While being completely equivalent to the original conditional Eq. (13), Eqs. (18) or (20) immediately suggest that the full treatment of nonadiabatic effects from some (approximate) reference ground-state-like conditional wavefunction will involve couplings (that manifest themselves as off-diagonal elements in the conditional Hamiltonian matrix of Eq. (18), for example) with the complementary space of conditional excited states. This qualitative picture will be turned into practical computational schemes in Sec. V, once the formalism has been merged (in the next section) with molecular KS-DFT and a reference (approximate) conditional wavefunction has been identified (See Sec. III.2).
@@ II.2 Molecular KS-DFT
respectively Fromager and Lasorne (2024). In the above construction, we have implicitly assumed that the nuclear and electronic density mappings of Eq. (25) can be achieved onto a pure electronically non-interacting molecular wavefunction ($\Psi^{\rm KS}$), by analogy with regular electronic KS-DFT. We shall refer to this assumption as molecular non-interacting $V$-representability. While the non-interacting $v$-representability is now better understood in the context of electronic KS-DFT Gonis (2019), the problem has not been addressed yet in the case of molecular KS-DFT. This is left for future work. Note that, for practical purposes, standard (purely electronic) density-functional approximations (DFAs) can be recycled within the adiabatic approximation that was deduced from an exact molecular generalization of the adiabatic connection formalism in Ref. Fromager and Lasorne (2024). This showed that the molecular Hxc functional can be written as [see Sec. IV of Ref Fromager and Lasorne (2024)]:
$\mathcal{E}_{\rm Hxc}[\Gamma,n]=\int dR\;\Gamma(R)\,\int_{0}^{1}d\lambda\,\langle\hat{W}_{\rm ee}\rangle_{\phi_{R}^{\lambda}[\Gamma,n]},$ (31)
with $\lambda$ being the electron-electron interaction strength scaling parameter and $\phi_{R}^{\lambda}[\Gamma,n]:r\mapsto{\Psi^{\lambda}[\Gamma,n](R,r)}/{\sqrt{\Gamma(R)}}$ the effective (conditional) electronic wavefunction of the partially-interacting molecule. Assuming that, for all $R$, $\phi_{R}^{\lambda}[\Gamma,n](r)$ matches its local-in-$R$ approximation $\phi^{\lambda}[n_{R}](r)$, that is the ground partially-interacting BO electronic state reproducing the electronic density $n_{R}:\mathbf{r}\mapsto n(R,\mathbf{r})$ at each nuclear geometry $R$, one can leverage the KS-DFT adiabatic connection formula,
$E_{\rm Hxc}[n]=\int_{0}^{1}d\lambda\,\langle\hat{W}_{\rm ee}\rangle_{\phi^{\lambda}[n]},$ (32)
@@
In this approximate picture, the electronic density $n_{R}$ is thought as being parameterized by the geometry $R$, hence the name “adiabatic” given to the approximation Fromager and Lasorne (2024). Its limitations and remedies will be discussed in a forthcoming paper (see also Refs. 23; 45; 43). Later in Sec. IV, the theory will be applied to a model system for which numerically exact densities can be evaluated and the corresponding KS potentials can be determined by reverse-engineering.
@@ III.2 Approximate KS-like one-electron conditional equation
thus isolating second-order geometrical derivatives (that is derivatives with respect to nuclear coordinates, the latter defining what is often called the molecular geometry) in the second term on the right-hand side. The existence of the latter term echoes the usual considerations related to the account (or not) of diagonal BO corrections (DBOCs) and reduced mass corrections within a BH picture Scherrer et al. (2017); Mátyus and Teufel (2019); Maskri and Joubert-Doriol (2022), which are known to be usually small (perturbative) around equilibrium geometries (far from crossings) and often neglected within the BO approximation. While it might be treated within perturbation theory Li et al. (2025); Tu and Gross (2025b); Schild et al. (2016), the above matrix form indicating clearly and explicitly how the (to-be-defined) unperturbed conditional solutions would couple, a more involved diagonalization approach can also be formulated, as shown later in Sec. V. In this section, we will simply neglect second-order geometrical derivatives, thus leading to the approximate conditional KS equation,
@@
where the index $\mu\geq 0$ has been introduced for labelling a specific solution. The ground-state ($\mu=0$) solution is an approximation to the conditional KS electronic wavefunction $\Phi^{\rm KS}_{R}(r)$. As discussed in Sec. V, excited-state ($\mu>0$) solutions to this problem can be used as a basis for approaching $\Phi^{\rm KS}_{R}(r)$ even further. In the context of perturbation theory, they would play the role of the perturbers. Let us remark that the above Hamiltonian is not Hermitian, and its eigenstates are not orthogonal. This feature can be made more explicit by solving Eq. (49) in perturbation theory, as shown in detail in Appendix B (see, in particular, the overlap expression through first order in Eq. (104)).
@@
according to the definition of $\hat{h}^{(0)}(R)$ (see Eqs. (37) and (39d)). Note that, in practice, the zeroth-order coupling potential ${u}^{\rm coup(0)}_{\rm ne}(R)$ would be evaluated from the ground-state ($\mu=0$) solution via the KS vector potential expression (see Eq. (38c)) combined with Eqs. (15) and (39b). We also note in passing that, despite the non-Hermiticity of the Hamiltonian in Eq. (50), the $\mu$-dependent (i.e., solution-dependent) contribution to the energy (the sum on the right-hand side of Eq. (52)) remains real-valued as long as the deviation from Hermiticity can be treated as a perturbation, as readily seen from Eqs. (100) and (106) [see Appendix B for further details]. Stronger coupling regimes would require further investigation. This is left for future work.
@@
Let us finally stress that Eq. (50) is more advanced than the beyond-BO KS Eq. (40) of Ref. 14, because of the additional first-order geometrical derivative contribution ${u}^{\rm coup(1)}_{\rm ne}(R)\partial/\partial R$. It is also substantially different from the exactly-factorized DFT of Wang et al. Wang et al. (2025) (see also Ref. 43), where the beyond-BO KS potential operator remains multiplicative (see Eq. (17) in Ref. 45).
@@ IV Model system study
The BO ground and first excited electronic PESs are drawn on top panel of Fig 1 together with the exact ground state nuclear density as a function of $R$. The latter was obtained by the means of a Discrete Variable Representation (DVR) Colbert and Miller (1992) calculation of the molecular ground-state wavefunction. On the bottom panel, we compare the electronic density difference between atomic sites obtained in the BO approximation to the exact result. The BO solution undergoes a sharp charge transfer at the avoided crossing between $E_{1}$ and $E_{2}$, going from a strongly ionic character to an almost symmetrical split of electronic density on each atom. In contrast, the exact solution shows taking non-adiabatic effects into account displaces the distance of equilibrium between ionic and neutral character by 1 atomic unit and makes the transition much smoother. The model presents strong non-adiabatic effects while being simple enough so that we can reverse-engineer the KS potentials reproducing the exact (beyond-BO) nuclear and electronic densities. We leave the detailed exposition of our parametrization and numerical details to a follow-up publication dedicated to approximating the Kohn-Sham (KS) energy functional. Presently, we focus on developing practical approximations to the conditional KS wavefunction equation (41) with the nuclear-nuclear and nuclear-electron (simply referred to as electronic in the following) KS potentials, be it exact or approximate, being given. We emphasize this is an entirely separate (yet significant) matter, and we treat it as such by using the same KS potentials in the full solution to Eq. 40, with a full account of second-order geometric derivatives, and its first order approximation Eq. (48).
@@
On Fig. 2 we plot the geometric first derivative of the exact KS coefficients in the singlet basis ($\Phi_{R}^{\rm KS}=\sum^{3}_{i=1}C^{\rm KS}_{i}(R)\;\Phi_{i}$). Those were obtained by solving the KS analog to molecular Schrödinger Eq. (1) in full (full lines), with exact nuclear density and electronic KS potential as input. As a comparison, we computed the prediction that Eq. (48) [which is equivalent to the one-electron conditional KS Eq. (50)] gives for the coefficients’ first derivative when fed with exact KS coefficients, electronic potential and nuclear density (dashed lines). The procedure is outlined in Appendix C. The singular behavior around $R=4$ a.u. corresponds to the maximum of the ground-state nuclear density and thus simply reflects that when ${u}^{\rm coup(1)}_{\rm ne}(R)$ vanishes, Eq. (48) does not tell us any information about the change of the electronic wavefunction at this position. There, it simplifies to a (local) eigenvalue problem with a KS electronic potential imbuing non-local (beyond-BO) effects. Thus, would the approximate solution be satisfactory everywhere else, no pathologic behavior would actually emerge. On that point, the prediction of Eq. (48) is reasonably close to the exact one so that our proposed decomposition into a ”first order” solution (Eq. (50)) followed by a perturbative incorporation of second-order derivative contributions holds promise.
@@ V Outlook: Correlating the conditional KS wavefunction
$[{\bf S}(R)]_{\nu\mu}=\int dr\,\left[\Phi^{\nu}_{R}(r)\right]^{*}\Phi^{\mu}_{R}(r)=\langle\Phi^{\nu}_{R}|\Phi^{\mu}_{R}\rangle,$ (60)
respectively. The non-orthogonality of the Slater determinants stems from the non-Hermitian nature of the “first-order” approximate Hamiltonian in Eq. (49). As mentioned previously, this can be made more explicit when treating the deviation from Hermiticity in perturbation theory (see Appendix B for further details). In fact, as readily seen from Eq. (104), the overlap matrix elements are (through first order in perturbation theory) proportional to the non-adiabatic couplings between the unperturbed solutions, i.e., the solutions to Eq. (49) where the first-order coupling term ${u}^{\rm coup(1)}_{\rm ne}(R)$ has been neglected. In general, a non-Hermitian operator has complex eigenvalues, but cases where they happen to be real exist Surján et al. (2024), as in the perturbative regime described in Appendix B (see, in particular, the non-vanishing second-order correction to the energy in Eq. (106)), due to the nature of the terms involved in assembling their expressions.
A working equation can finally be obtained through multiplication by the marginal nuclear density and integration over the geometry, thus leading to
@@
As a final note, let us point out that a possibly more accurate one-electron orbital approximation to Eq. (41) would consist in only neglecting second-order derivative terms involving two different orbitals of the conditional wavefunction. Indeed, by writing
$\begin{split}\frac{\partial^{2}\Phi^{\mu}_{R}}{\partial R^{2}}=&\sum^{N_{\rm e}}_{i=1}\left|\varphi_{R}^{(1^{\mu})}\ldots\dfrac{\partial^{2}\varphi_{R}^{(i^{\mu})}}{\partial R^{2}}\ldots\varphi_{R}^{(N_{\rm e}^{\mu})}\right|\\ &+\sum_{i\neq j}^{N_{e}}\left|\varphi_{R}^{(1^{\mu})}\ldots\dfrac{\partial\varphi_{R}^{(i^{\mu})}}{\partial R}\ldots\dfrac{\partial\varphi_{R}^{(j^{\mu})}}{\partial R}\ldots\varphi_{R}^{(N_{\rm e}^{\mu})}\right|,\end{split}$ (66)
we see that neglecting the cross derivative terms (on the second line of the above equation) in the evaluation of second-order derivatives in Eq. (41) allows us to recover a one-electron-like picture with the following (still approximate) conditional KS orbital equation,
@@
This motivates the separation of second-order geometrical derivative terms into a one-electron component (first line of Eq. (66)),
$\frac{\partial^{(1)}}{\partial R}\dfrac{\partial\tilde{\Phi}^{\mu}_{R}}{\partial R}\equiv\sum^{N_{\rm e}}_{i=1}\left|\tilde{\varphi}_{R}^{(1^{\mu})}\ldots\dfrac{\partial^{2}\tilde{\varphi}_{R}^{(i^{\mu})}}{\partial R^{2}}\ldots\tilde{\varphi}_{R}^{(N_{\rm e}^{\mu})}\right|,$ (68)
and the cross-derivative terms as the remainder,
$\frac{\partial^{(2)}}{\partial R}\frac{\partial\tilde{\Phi}^{\mu}_{R}(r)}{\partial R}=\dfrac{\partial^{2}\tilde{\Phi}^{\mu}_{R}(r)}{\partial R^{2}}-\frac{\partial^{(1)}}{\partial R}\dfrac{\partial\tilde{\Phi}^{\mu}_{R}}{\partial R}.$ (69)
The above contribution describes the first-order geometrical derivative of two different orbitals within the Slater determinant $\tilde{\Phi}^{\mu}_{R}{\equiv}\left|\tilde{\varphi}_{R}^{(1^{\mu})}\tilde{\varphi}_{R}^{(2^{\mu})}\ldots\tilde{\varphi}_{R}^{(N_{\rm e}^{\mu})}\right|$ and does model, as such, correlation effects. The set of Slater determinants generated from these orbitals fulfills the conditional KS equation
@@
which is expected to be more accurate than Eq. (48), while still being amenable to a one-electron computational treatment. In this case, the total KS energies read
@@
Finally comes the question of how to generate the in-principle exact correlated conditional KS wavefunction $\Phi^{\rm KS}_{R}$ from the alternative basis of Slater determinants $\{\tilde{\Phi}^{\mu}_{R}\}$. In fact, we can simply proceed by analogy with Eq. (56). Indeed, from the following expansion,
@@
and Eq. (70), the exact conditional KS Eq. (40) becomes, according to Eq. (69),
$\begin{split}&\sum_{\mu\geq 0}\tilde{C}_{\mu}\left[\begin{matrix}\tilde{\mathcal{E}}^{\mu}(R)&-\frac{1}{2M}\frac{\partial^{(2)}}{\partial R}\\ 0&1\end{matrix}\right]\left[\begin{matrix}\tilde{\Phi}^{\mu}_{R}(r)\\ \frac{\partial\tilde{\Phi}^{\mu}_{R}(r)}{\partial R}\end{matrix}\right]\\ &=\left[\begin{matrix}\mathcal{E}^{\rm KS}(R)&0\\ 0&1\end{matrix}\right]\sum_{\mu\geq 0}\tilde{C}_{\mu}\left[\begin{matrix}\tilde{\Phi}^{\mu}_{R}(r)\\ \frac{\partial\tilde{\Phi}^{\mu}_{R}(r)}{\partial R}\end{matrix}\right],\,\forall R,\,\forall r,\end{split}$ (73)
or, equivalently,
$\begin{split}&\sum_{\mu\geq 0}\tilde{C}_{\mu}\left[\begin{matrix}\tilde{\mathcal{E}}^{\mu}(R)&-\frac{1}{2M}\frac{\partial^{(2)}}{\partial R}\end{matrix}\right]\left[\begin{matrix}\tilde{\Phi}^{\mu}_{R}(r)\\ \frac{\partial\tilde{\Phi}^{\mu}_{R}(r)}{\partial R}\end{matrix}\right]\\ &=\mathcal{E}^{\rm KS}(R)\sum_{\mu\geq 0}\tilde{C}_{\mu}\tilde{\Phi}^{\mu}_{R}(r),\,\forall R,\,\forall r,\end{split}$ (74)
which can be solved along the same lines as before (see Eqs. (58)-(65)).